Post Corporation paid $330,000 cash for 75% of the outstanding common stock of Soft Company on January 1, 2020. There was no control premium and the fair value of the noncontrolling interest was $110,000 on January 1, 2020. Differences between book value and fair value of the net identifiable assets of Soft Company on January 1, 2020, were limited to the following:

                                                                      Book value        Fair value         

            Inventories                                          $ 25,000          $ 35,000                       

            Building (net)                                       175,000            173,000                       

Required:

(i) Prepare the working paper elimination entries E and R (in journal entry format) for Post Corporation and subsidiary on January 1, 2020.                                                

(ii) Complete the following working paper:

Working paper for consolidated balance sheet on date of business combination, January 1, 2020                      

Post

Dr (Cr)

Soft

Dr (Cr)

Adjustments & Eliminations

Consolidated

Dr (Cr)

Debits

Credits

Cash

20,000

30,000

Inventories

170,000

25,000

Investment in Soft

330,000

Building (net)

230,000

175,000

Accounts payable

(120,000)

(110,000)

Common stock

(380,000)

(25,000)

Add. paid-in capital

(100,000)

(30,000)

Retained earnings

(150,000)

(65,000)

       Total

          0

           0
Answer #1

1.

Working paper elimination entries E and R (in journal entry format) for Post Corporation and subsidiary on January 1, 2020.
Sl.No. Name of the Account Debit Credit
1 Inventory Account 5000
To Fair Value Adjustment Account 5000
(being adjustment made for the difference)
2 Fair value adjustment Account 2000
To Buildings Account 2000
(being adjustment made for the difference)
3 Equity 92250
Good Will Account 237750
To Investment in Soft 330000
(being the equity accounts eliminated)
4 Equity 30750
To Non Controlling interest 30750
(being the equity accounts eliminated)

2. Working paper for consolidated balance sheet on date of business combination, January 1,2020

Particulars Post
Dr (cr)
Soft
Dr (cr)
Adjustments & Eliminations
Dr (cr)
Consolidated
Dr (cr)
Cash 20,000 30,000 50,000
Inventories 170,000 25,000 5,000 200,000
Investment in Soft 330,000 (330,000)
Building (net) 230,000 175,000 (2,000) 403,000
Good Will 237,750 237,750
Accounts payable (120,000) (110,000) (230,000)
Common stock (380,000) (25,000) 25,000 (380,000)
Add. paid-in capital (100,000) (30,000) 30,000 (100,000)
Retained earnings (150,000) (65,000) 65,000 (150,000)
Non-Controlling interest (30,750) (30,750)
       Total           0            0 0 0

Calculation:

Net Assets Value of Soft Company:

Total Assets - 230,000

Fair Value Adjustments - 3,000

(5000-2000)

Total Assets - 233,000

Less: Liabilities - 110,000

Net Assets value - 123,000

Share of Post Corporation = 123,000 * 75% = 92,250

Share of Minority Interest Shareholders = 123,000*25% = 30,750